A monopolist can sell 300 units of output for $45 per unit. Alternatively, it can sell 301 units of output for $44.60 per unit. The marginal revenue of the 301st unit of output is? a) -$0.40. b) -$75.40. C)$75.40. d) -$120.00.
Answer #1

Marginal revenue is change in revenue due to 1 unit change in output that means to calculate marginal revenue we have to calculate the total revenue change due to produce additional output.

Now total revenue = Price \times Quantity

therefore total revenue for 300 unit is 45 \times 300 = 13500

And total revenue for 301 unit is 44.60 \times 301 = 13424.6

Now

Marginal revenue = total revenue = (13424.6 - 13500) = -75.40

That means option B is correct